OpenStudy (anonymous):

solve in scientific notation: 66,000/0.2

OpenStudy (akashdeepdeb):

\[\frac{ 60000 }{ 0.2 } = \frac{ 600000 }{ 2 } = 300000 = 3 * 10^5\] JUST SIMPLIFY IT TILL THE END! And make sure it is in the form = a * 10^n [Where a is gretaer than or equal to 1 and less than 10] Got it? :)

OpenStudy (akashdeepdeb):

@meandpuppies ?

OpenStudy (anonymous):

Ok, 3.3x10^5. I didn't think about doing it that way, but makes sense:)

OpenStudy (phi):

solve in scientific notation: 66,000/0.2 change 66,000 to scientific notation change 0.2 to scientific notation